OpenStudy (maheshmeghwal9):

The two ends of a train moving with uniform acceleration pass a certain point with velocities "u" & "v". The velocity with which the middle portion of the train passes the same point is -???????????????????????????????????????????????????????????????????????????????????????????????? Plz help:)

OpenStudy (unklerhaukus):

\[\frac {u+v}{2}\]?

OpenStudy (maheshmeghwal9):

Its answer is \[\LARGE \color{red}{\sqrt{\frac{u^2+v^2}{2}.}}\]

OpenStudy (unklerhaukus):

hm.

OpenStudy (maheshmeghwal9):

hey @mathslover plz help kar yar:)

mathslover (mathslover):

wait

OpenStudy (maheshmeghwal9):

Hmmmmmmmmmmmmmm..................................... KKKKKKKKKKKKKKKKKKKKKKKKKKKKKKKK!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (maheshmeghwal9):

@A.Avinash_Goutham @KingGeorge Plz help:)

mathslover (mathslover):

@maheshmeghwal9 \[v^2=u^2+2as\] \[as=\frac{v^2-u^2}{2}\] let x be the velocity of the middle part of train.. \[x^2=u^2+as\] \[x=\sqrt{u^2+as}\] \[x=\sqrt{u^2+\frac{v^2-u^2}{2}}\] \[x=\sqrt{\frac{v^2+u^2}{2}}\]

mathslover (mathslover):

got it ?

mathslover (mathslover):

have to go thanks ..

OpenStudy (maheshmeghwal9):

no I didn't gt it:/

OpenStudy (maheshmeghwal9):

how did u write x^2=u^2+as

OpenStudy (maheshmeghwal9):

I men why?

OpenStudy (maheshmeghwal9):

mean*

mathslover (mathslover):

x^2 = u^2 + 2*a*s/2 ... right ?

mathslover (mathslover):

s/2 = half of length of train ...

mathslover (mathslover):

understood now ?

OpenStudy (maheshmeghwal9):

Why did u apply this formula? v^2=u^2+2as.

OpenStudy (maheshmeghwal9):

& I think "s" is distance. How can it be length of the train?

OpenStudy (maheshmeghwal9):

@mathslover

mathslover (mathslover):

does this make clear to you ? @maheshmeghwal9

OpenStudy (maheshmeghwal9):

very much clear dude. thanx a lot:)

mathslover (mathslover):

my pleasure @maheshmeghwal9

OpenStudy (maheshmeghwal9):

^_^ @mathslover

mathslover (mathslover):

u can close this question now..

OpenStudy (maheshmeghwal9):

ya I forgot that